# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v2

@@ 3.2 Identification of the shocks by non-Gaussianity
    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, $i=1,...,d$, has zero mean and unit variance.
@@
Note that the assumption of a unit variance of $e_{it}$, $i=1,...,d$, is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$. Also, due to the time-variation of the impact matrix (3.3), Assumption 1 does not imply that the reduced form innovations $u_{t}$ are independent nor that they are identically distributed.
The next lemma then shows that, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns at each $t$. This means that only the impact matrices that are obtained by reordering or changing the signs of the columns of $B_{y,t}$ are observationally equivalent. This lemma, which is mostly similar to Proposition 1 in Lanne et al. (2017), is used in the following sections to show that the parameters in the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified.
@@
We have concluded that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1. In other words, at each $t$‚ changing the ordering and signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to a model that is not observationally equivalent. The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1. If the impact matrix is time-invariant as in Lanne et al. (2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows that the structural shocks are identified up to ordering and signs. Such statistically identified shocks can then be labelled as economic shocks based on external information (see Lanne et al., 2017). However, when the impact matrix varies in time, two complications arise. First, the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, and it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified. Second, since the identification of $B_{y,t}$ at each $t$ is only up to ordering and signs of its columns, its unique identification requires that the parameters in the functional form of $B_{y,t}$ are restricted to fix the ordering and signs of the columns of $B_{y,t}$.
To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$. In order to fix the ordering and signs of the columns of $B_{y,t}$ it then suffices to fix the ordering and signs of the columns of $B_{1},...,B_{M}$. This requires such constraints to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{m}$ are not observationally equivalent at some $t$.
It is not immediately obvious what is the best strategy for fixing the ordering and signs of the columns of $B_{1},...,B_{M}$, as depending of the transition weights and distributions of the shocks, constraints that are not restrictive in some specifications can be overidentifying in others. Also, constraints that are identifying in some specifications may not be sufficient for identification in others. Therefore, it is useful to divide the discussion on identification into the cases of binary (i.e., $\alpha_{m,t}\in\{0,1\}$) and more general transition weights.
@@ 3.3 Identification of the shocks in structural TVAR models
We consider structural threshold VAR (TVAR) models here, as they are special cases of STVAR models, but the introduced identification results also directly extend to other models incorporating discrete regime switches, such as Markov-switching VAR models (Krolzig, 1997) and mixture VAR models (e.g., Fong et al., 2007). Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$. Then, at every $t$, the process is completely in one of the regimes and the impact matrix is $B_{y,t}=B_{m}$ for the regime $m$ with $\alpha_{m,t}=1$. That is, we have structural TVAR model. By Lemma 2, the impact matrix $B_{y,t}$ is thereby identified up to ordering and signs of its columns in each of the regimes, implying that identification is obtained by fixing the ordering and signs of the columns of $B_{1},...,B_{M}$. However, whether this is overidentifying depends on the distributions of the shocks $e_{t}=(e_{1t},...,e_{dt})$.
For instance, suppose the shocks follow independent Student’s $t$ distributions. Then, the signs of the columns of $B_{m}$ can be fixed in each regime separately without loss of generality due to the symmetry of the Student’s $t$ distribution. But fixing the ordering of the columns in each regime separately would generally be overidentifying because the $i$th column of $B_{m}$ is related to the $i$th shock, and each shock has its own degrees-of-freedom parameter that affects the density of the distribution (and is in our specification shared across the regimes). Fixing the ordering of the columns of $B_{1}$ would thereby effectively fix the ordering of the columns of $B_{2},...,B_{M}$ without restricting the model when the values of the degrees-of-freedom parameter are not equal for any of the shocks. On the other hand, if the degrees-of-freedom parameter values of some of the shocks are equal, the ordering of the corresponding columns can be changed in each regime separately without changing the underlying model. If the degrees of freedom parameter values are close to each other, reordering the columns of $B_{m}$ in one of the regimes without reordering them in the other regimes accordingly would have only a small effect of the density of the model (or to the fit of an estimated model), thereby resulting in weak identification with respect to the ordering of the columns of $B_{2},...,B_{M}$.
As the required identification conditions depend on the distributions of the shocks, it useful to make some assumptions on them. Namely, we assume that the distributions of the shocks are continuous and none of the them are identical to each other in the sense that the values of their density functions are unequal almost everywhere in $\mathbb{R}$. This assumption allows to effectively fix the ordering of the columns of $B_{2},...,B_{M}$ by fixing the ordering of the columns of $B_{1}$, since changing the ordering would change the distribution of the corresponding shocks in some of the regimes relative to Regime 1. If this assumption is close to break (e.g., when the degrees-of-freedom parameter values of independent Student’s $t$ distributions are close to each other or very large), identification is weak with respect to the ordering of the columns. Labelling the shocks based on the estimates of the columns of each $B_{m}$ can then be unreliable, as small changes in the data could result in changes in the ordering of columns in some of the regimes. Therefore, we propose a procedure to facilitate reliable labelling of the shocks in the presence of such weak identification in Section 3.5.
The following proposition, which is proved in Appendix B.3, formally states the identification results for TVAR models.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible, Assumption 1 satisfied and $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$. Denote the density function of the distribution of the shock $e_{it}$ as $d_{e_{i}}(\cdot;0,1,\nu_{i})$, where $\nu_{i}$ denotes the parameters of the distribution other than mean and variance. Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$ and assume the following conditions hold:
@@
    the signs of the columns of $B_{m}$ are fixed for each $m=1,...,M$,
@@
    $e_{it}$ follow continuous distributions with $d_{e_{i}}(x;0,1,\nu_{i})\neq d_{e_{j}}(x;0,1,\nu_{j})$ almost everywhere in $x\in\mathbb{R}$ for $i\neq j=1,....,d$.
Then, almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{T}$, the matrices $B_{1},...,B_{M}$ are uniquely identified.
Condition (a) of Proposition 1 grants arbitrary identification with respect to the signs of the shocks. If the distributions of the shocks are symmetric, which is typically assumed in nonlinear SVAR models, any fixed signs can be assumed without loss of generality. However, if the distributions of the shocks are not symmetric, the signs of the columns of the impact matrices may affect the density of the model, implying that fixing the signs is an overidentifying restriction. Condition (b) fixes the ordering of the columns of $B_{1}$. Any fixed ordering of the columns can be assumed without loss generality. For instance, if the first nonzero entry in each column of $B_{1}$ is assumed positive, Condition (a) is satisfied by the positivity assumption, and Condition (b) is satisfied by assuming that these entries are in a decreasing order, given that none of the entries are exactly equal to each other.
Condition (c) is concerned with the distributions of the shocks, and it is satisfied by many of the popular distributions such as Student’s $t$ distributions with different values of the degrees-of-freedom parameters. It ensures that changing the ordering of the columns of $B_{2},...,B_{M}$ while keeping the ordering of the columns of $B_{1}$ fixed would change the joint density of the model. Thus, combined with Condition (b), it leads to identification with respect to ordering of the columns of $B_{1},...,B_{M}$. Finally, the result of Proposition 1 holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{T}$, because in some special cases (e.g., if all observations are zero), changing the ordering of the columns of $B_{2},...,B_{M}$ would not change the joint density of the model. Since the probability of obtaining such special cases is zero, this does not have any practical implications.
@@ 3.4 Identification of the shocks in structural STVAR models
Having established the identification result for structural TVAR models, or more generally to models incorporating discrete regime switches, we now assume that the transition weights exhibit more variation. Because the variation in the transition weights then translates to variation in the impact matrix (3.3), which is at each $t$ identified up to ordering and signs of its columns (by Lemma 2), the identification result can be stated without making further assumptions about the distributions of the shocks other than those in Assumption 1. Specifically, under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied. Suppose the following conditions hold:
@@
    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors $(\alpha_{n_{1},t},...,\alpha_{n_{M^{\circ}},t})$, $n_{l}\in\mathbb{M^{\circ}}$, $\mathbb{M^{\circ}}\subset\{1,...,M\}$, where $M^{\circ}\geq 2$ denotes the number of elements in $\mathbb{M^{\circ}}$, take at least $M+1$ linearly independent values.
Then, almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, the matrices $B_{1},...,B_{M}$ are uniquely identified.
@@
Proposition 2 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified. The result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that there can exist some special cases of $B_{1},...,B_{M}$ such that fixing the ordering and signs of the columns of $B_{1}$ does not fix the ordering and signs of the columns of $B_{2},...,B_{M}$, but this set has a Lebesgue measure zero in $\mathbb{R}^{d\times dM}$.
Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$. Switching the ordering of the columns of $B_{y,t}$ merely equals switching the ordering of the shocks, so any fixed ordering of the columns of $B_{1}$ can be assumed without loss of generality. Furthermore, if the distributions of the shocks are symmetric, which is typically assumed in nonlinear SVAR models, also the signs of the columns of $B_{1}$ can be fixed without loss of generality, as switching the sign of a column of $B_{y,t}$ equals switching the sign of the corresponding shock. The ordering and signs of the columns of $B_{1}$ can be fixed, for example, by assuming that the first nonzero element in each column of $B_{1}$ is positive, these elements are not exactly equal to each other, and they are in a decreasing ordering. The assumption of unequal first nonzero elemenents is not restrictive in practice, as the probability of obtaining an estimate such that some of the elements are exactly equal to each other within the numerical accuracy of modern statistical software is extremely small. See Lanne et al. (2017) for a more complicated strategy for fixing the ordering and signs (or more generally scalings) of the columns of a static impact matrix that is valid without such an additional assumption.
@@
Condition (B) of Proposition 2 states that the transition weights should, for some $t$, be strictly positive for all of the regimes. If there are only two regimes, the requirement is merely that the regime-switches are not always discrete (in which case Proposition 1 would be more relevant). Condition (C), in turn, states that the transition weights should exhibit at least a certain small degree of variation. For instance, if there are only two regimes in the model, the transition weights should take at least $M+1$ distinct values, as the vectors $(\alpha_{1,t},\alpha_{2,t})$ with distinct $\alpha_{1,t}$ are always linearly independent due to the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$. In practice, Condition (C) should be satisfied for the observed time period $t=1,...,T$. This assumption is not, however, very restrictive, as the transition weights vary either depending on the preceding observations or some exogenous variables, and it is very uncommon that their variation would result in only a few distinct values of the transition weights (except when the regime switches are discrete).
The concerns of weak identification with respect to the ordering (and signs) of the columns of $B_{2},...,B_{M}$ with binary transition weights, discussed in Section 3.3, also translate to smooth transition models when the estimates of the transition weights are such that the weights are mostly close to zero or one. Surprisingly, our experience shows that even when there is a high degree of variation in the transition weights, the identification with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$ often (if not virtually always) appears to be very weak in practice, at least if the parameters are estimated by the method of maximum likelihood. Specifically, there are frequently a number of local maxima of the log-likelihood function with log-likelihoods very close to each other, corresponding to mostly very similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$.
Weak identification with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$ can be a problem in practice for labelling the shocks, as a given column in $B_{y,t}$ need not correspond to the same shock for all $t=1,...,T$. In this case, the (generalized) impulse response functions would not estimate of the effects of the intended shock but of a linear combination of the structural shocks. As a solution, in the next section, we propose a procedure that utilizes the almost equally well fitting local solutions obtainable by our estimation method (discussed in Section 4). If necessary, the identification can also be strengthened by imposing economic overidentifying short-run restrictions, which are testable.
3.5 Labelling the shocks

### v3

@@ 3.2 Identification of the shocks by non-Gaussianity
    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, $i=1,...,d$, follows a continuous distribution with zero mean and unit variance.
@@
Note that the assumption of a unit variance is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$. Also, due to the time-variation of the impact matrix (3.3), Assumption 1 does not imply that the reduced form innovations $u_{t}$ are independent nor that they are identically distributed.
The next lemma then shows that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns. This means that only the impact matrices that are obtained by reordering or changing the signs of the columns of $B_{y,t}$ are observationally equivalent. This lemma, which is partly similar to Proposition 1 in Lanne et al. (2017), is used in the following sections to show that the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified. Moreover, Lemma 2 also shows that if the transition weights are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic (2.4) process, then also the autoregressive (AR) as well as the weight function parameters are identified. Specific weight functions are assumed here for simplicity, but the result can be extended to other suitable weight functions as well.
@@
Moreover, suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$. If the following conditions hold:
  1. (1)
    the transition weights $\alpha_{m,t}$, $m=1,...,M$, are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic process (2.4) (with $M=2$ in the latter), and
  2. (2)
    $(\phi_{m,0},\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))\neq(\phi_{n,0},\text{vec}(A_{n,1}),...,\text{vec}(A_{n,p}))$ for all $m\neq n\in\{1,...,M\}$,
the parameters $\phi_{m,0},A_{m,1},...A_{m,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. For threshold models (2.3), also the parameters $r_{1},...,r_{m-1}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. For logistic models (2.4), the parameters $c,\gamma$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, where $z_{t}$ is the switching variable defined in (2.4).
@@
We have concluded that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1. In other words, at each $t$‚ changing the ordering or signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to an observationally distinct model. The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1. Identification of the AR and weight function parameters, in turn, is achieved only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, because in some special cases of degenerate data (e.g., if the regime does not switch), there is insufficient variation to distinguish the parameters.
If the impact matrix is time-invariant as in Lanne et al. (2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows from Lemma 2 that the structural shocks are identified up to ordering and signs. Such statistically identified shocks can then often be labelled as economic shocks based on external information (see Lanne et al., 2017). However, when the impact matrix varies in time, two complications arise. First, the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, and it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified. Second, since the identification of $B_{y,t}$ at each $t$ is only up to ordering and signs of its columns, its unique identification requires that the parameters in the functional form of $B_{y,t}$ are restricted to fix the ordering and signs of the columns of $B_{y,t}$.
To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$. To fix the ordering and signs of the columns of $B_{y,t}$, it suffices to fix the ordering and signs of the columns of $B_{1},...,B_{M}$. This requires such constraints to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{m}$ are not observationally equivalent at some $t$.
It is not immediately obvious what is the best strategy for fixing the ordering and signs of the columns of $B_{1},...,B_{M}$, as depending on the transition weights and distributions of the shocks, constraints that are not restrictive in some specifications can be overidentifying in others. Also, constraints that are identifying in some specifications may not be sufficient for identification in others. Therefore, it is useful to divide the discussion on identification into the cases of binary (i.e., $\alpha_{m,t}\in\{0,1\}$) and more general transition weights.
@@ 3.3 Identification of the shocks in structural TVAR models
Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$. Then, at every $t$, the process is completely in one of the regimes and the impact matrix is $B_{y,t}=B_{m}$ for the regime $m$ with $\alpha_{m,t}=1$. That is, we have a structural TVAR model. By Lemma 2, the impact matrix $B_{y,t}$ is thereby identified up to ordering and signs of its columns in each of the regimes, implying that identification is obtained with constraints that fix the ordering and signs of the columns the impact matrix in each regime. To state the identification result, a specific shock distribution is assumed here for simplicity. Namely, we assume that each shock $e_{it}$, $i=1,...,d$, follows the skewed $t$-distribution of Hansen (1994), which can flexibly capture fat tails and skewness in the data, nests the Gaussian distribution as a special case, and is discussed in detail in Section 4.111The identification result can be extended to other suitable shock distributions using arguments similar to those in the proof of Proposition 1 in Appendix B.3.
The following proposition, which is proved in Appendix B.3, states the identification result for the TVAR models.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible, Assumption 1 satisfied, $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$, and the shocks $e_{it}$, $i=1,...,d$, following the skewed $t$-distribution defined in Equation (4.1) with the parameters $\nu_{i}>2$ and $\lambda_{i}\in(-1,1)$ controlling kurtosis and skewness, respectively. Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
@@
    the ordering and signs of the columns of $B_{1}$ are fixed,
@@
    $\lambda_{i}\neq 0$ for all $i=1,...,d$.
Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
Condition (a) of Proposition 1 fixes the ordering and signs of $B_{1}$, which effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ under Conditions (b) and (c), respectively. Any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality (e.g., by assuming that the first nonzero entry in each column is positive and the first nonzero entries are in a decreasing order, given that none of them are equal). If Condition (b) fails for some $i,j\in\{1,...,d\}$, the distributions of these shocks are identical, implying that fixing the ordering of the columns of $B_{1}$ does not fix the ordering corresponding of the columns of $B_{2},...,B_{M}$. In such a case, the ordering of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality, yielding identification. If Condition (c), in turn, fails for some $i\in\{1,...,d\}$, the signs of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality, yielding identification. The result of Proposition 1 holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, because in some special cases (e.g., if all observations are zero), changing the ordering of the columns of $B_{2},...,B_{M}$ does not change the joint density of the model.
If Condition (b) or (c) is close to break, the identification is weak with respect to the ordering or signs of the corresponding columns of $B_{2},...,B_{M}$, respectively. Somewhat surprisingly, our experience shows that even when Conditions (b) and (c) are not close to break, the identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, at least if the parameters are estimated by the method of maximum likelihood or penalized maximum likelihood (discussed in Section 4). Specifically, there are frequently a number of local maxima of the (penalized) log-likelihood function with (penalized) log-likelihoods close to each other, corresponding to otherwise largely similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$. As discussed in Section 3.5, this weak identification should be appropriately dealt with when labelling the shocks, for instance, by combining Proposition 1 with additional identifying information.
@@ 3.4 Identification of the shocks in structural STVAR models
Having established the identification result for structural TVAR models, we now consider the transition weights that exhibit more variation. The transition weights are here assumed to be either exogenous (nonrandom) or logistic as defined in (2.4), so that the result of Lemma 2 establishing identification of the transition weights can be exploited. Because the variation in the transition weights translates to variation in the impact matrix (3.3), which is at each $t$ identified up to ordering and signs of its columns by Lemma 2, the identification result can be stated without making further assumptions about the distributions of the shocks other than those in Assumption 1. Specifically, under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied, and let the transition weights be either exogenous (nonrandom) or follow the logistic process defined in (2.4) (with $M=2$ assumed). Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
@@
    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors take at least $M+1$ linearly independent values.
Then, for models with logistic transition weights, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, and $z_{t}\in\mathbb{R}^{T}$, where $z_{t}$ is the switching variable defined in (2.4). For models with nonrandom exogenous transition weights, this result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
@@
Proposition 2 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified. The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as there may exist some special cases of $B_{1},...,B_{M}$ such that fixing the ordering and signs of the columns of $B_{1}$ does not fix the ordering and signs of the columns of $B_{2},...,B_{M}$, but this set has a Lebesgue measure zero in $\mathbb{R}^{d\times dM}$. For logistic models, the identification holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, as special cases of degenerate data may lack sufficient variation to uniquely identify the transition weights (see Lemma 2).
Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$. If the shocks follow the skewed $t$-distributions defined in (4.1), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality, as reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related skewness parameter. Condition (B) states that the transition weights should, for some $t$, be strictly positive for all of the regimes. If there are only two regimes, the requirement is merely that the regime-switches are not always discrete, in which case Proposition 1 would be more relevant.
@@
Condition (C) of Proposition 2 states that the transition weights should exhibit at least a certain small degree of variation. For instance, if there are only two regimes in the model, the transition weights should take at least $M+1$ distinct values, as the vectors $(\alpha_{1,t},\alpha_{2,t})$ with distinct $\alpha_{1,t}$ are always linearly independent due to the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$. This assumption is not, however, very restrictive, as the transition weights vary either depending on the preceding observations or some exogenous variables, and it is very uncommon that their variation would result in only a few distinct values of the transition weights (except when the regime switches are discrete). Similarly to the structural TVAR model discussed in Section 3.3, our experience shows that the identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately dealt with when labelling the shocks, as discussed in the next section.
3.5 Labelling the shocks and blended identification